Can you help me with this problem?
The velocity of an object that has mass a mass of 25kg is 15 m/s. It is accelerated by a force until it reaches a velocity of 40 m/s. This is accomplished in 5 seconds. What is the force that caused this acceleration?

Answer by scott8148(6628)

the acceleration is the change in velocity divided by the time (uniform acceleration is assumed)

a = (40 - 15) / 5 = 5 m/s^2

force equals the mass multiplied by the acceleration (Newton's 2nd law)

f = 25 * 5 = 125 N (guess what the N stands for)